# Reversibility in finite-dimensional collapse dynamics

### v3

@@ Abstract
We study finite dimensional quantum systems with arbitrary collapse events, establishing, under no-information-erasure conditions, a structural no-go for trajectory-level operational irreversibility. More precisely, we prove that, for every physically admissible selector of the collapse dynamics, there exists a topologically closed, forward-invariant subset of the projective state space on which any two states can be connected with arbitrarily fine Fubini–Study precision and arbitrarily small integrated energetic cost. This shows that the preservation of information along a realized branch of outcomes guarantees islands of quasi-reversibility, while genuine irreversibility requires additional ingredients such as non-compactness or information erasure.
@@ 1.1 Reversibility in unitary versus collapse dynamics
We remark that is not the standard stance in the quantum–trajectories literature: the seemingly intractable irregularity, that such state–wise selectors permit, is usually avoided, for instance by averaging over outcomes or retaining stochasticity, where continuity and stability are recovered in measure or in expectation rather than pointwise for a deterministic selector–induced map [9, 3, 2, 1, 21]. Our approach is instead to explore what can be said, from the point of view of recurrence, assuming only 1. and 2. above, so in particular allowing 3. We fix thus a realization map (the chosen selector of the dynamics) assuming only the (physics-dictated) properties 1-2 above, and allowing the density property (3), which implies everywhere discontinuity for the induced evolution map $T$.
@@ 1.2 The result
Once a realization map $D$ has been chosen, and an ambient Hamiltonian $H$ is fixed, with corresponding unitary group $U$, they determine a (very irregular) induced dynamics
@@
where points are unit vectors $u]\in\mathbb{P}(\mathcal{H})$.
Landauer’s principle (see [18]) bounds the minimal heat cost of logically irreversible operations (such as erasure), and conversely permits, in principle, arbitrarily low dissipation for logically reversible transformations in the quasi–static limit.
The choice of a selector for the dynamics can be interpreted as realizing the extreme, no-information-erasure limit of Landauer’s principle, in the sense highlighted by C.H. Bennett’s resolution of the Maxwell demon paradox [5, 6, 13]. In particular, we share the following position, expressed by Bennett (from [5]):
@@
Given two states $u],[v]\in\mathbb{P}(\mathcal{H})$, we can define the minimal forward cost of going from $[u]$ to $[v]$ under the dynamics law $T_{U,D}$ by
@@
We say that the dynamical law $T_{U,D}$ exhibits an operational arrow of time if there exist states $u],[v]\in\mathbb{P}(\mathcal{H})$ such that
@@
Conversely, on a subset $B\subset\mathbb{P}(\mathcal{H})$ we say that the operational arrow of time does not arise (or that $B$ is operationally reversible) if
@@
We can summarize our main result as follows.
Main result In the no–erasure regime, compactness enforces quasi–reversibility: given any compact quantum system and any realization of the collapse dynamics, there exists a topologically closed and invariant (with respect to the realization-induced evolution map) subsystem $S$ that is operationally reversible, meaning that, in $S$, quasi-returns with arbitrarily small Fubini-Study error and arbitrarily small energetic cost are allowed between any pair of states.
Genuine irreversibility, thus, requires additional ingredients: information loss/coarse–graining, non-compact limits, or coupling to reservoirs. From a conceptual perspective, this establishes a structural no-go: under finite-dimensional, no-erasure conditions, collapse alone cannot generate an operational arrow of time along individual realized branches.
@@ 1.3 Where a naive diagonal grid argument fails
It may look tempting to try to get a (naive) heuristic argument proving our main result as follows: fix $\epsilon>0$, cover $\mathbb{P}(\mathcal{H})$ by finitely many $d_{FS}$–balls ($\epsilon$–cells”), and note that along any long enough forward segment some cell must be visited twice; this yields an $\epsilon$–loop. Pushing $\epsilon\to 0$ seems to promise finer and finer quasi-loops. However, this argument is intrinsically scale–wise: as $\epsilon$ shrinks, the repeating cell and the base point inside it drift with $\epsilon$. Of course, by compactness the chosen cells admit a convergent subsequence, but in the absence of continuity the induced dynamics at the limit point is decoupled from the finite $\epsilon$–loops: $T$ at the limit may bear no relation to the images seen along the approximants. Consequently, the grid argument does not deliver a chain–recurrent point, let alone a topologically closed, invariant, internally strongly chain–transitive subsystem; nothing in this argument enforces nesting across scales of the pseudo–orbits.
@@
At the opposite extreme of the naive attempt just described, one could obtain an abstract pure existence result on the existence of a chain-transitive subsystem by using Zorn’s lemma on topologically closed, invariant subsets ordered by inclusion. Our aim, however, is to provide a constructive procedure, which is the only one, in our opinion, that works well with the operational character of the employed notion of reversibility. Therefore, not relying on the Axiom of Choice, we implement (in the Appendix) a transfinite argument that enforces both spatial nesting and dynamical coherence across scales. This procedure yields a strongly chain–recurrent point, and even a topologically closed TT–invariant subset that is internally $\mathcal{SC}_{d_{FS}}$–transitive (see Sections 2 and 3 for the relevant definitions). We remark that the two basic operations considered in the procedure, that is taking images of the map and passing to an accumulation point of an orbit, clearly have finite-precision analogues.
The paper is organized as follows: in Section 2 we introduce basic concepts and notation for the unitary evolution of the system and the abstract topological dynamics concepts that will be used later; in Section 3 we describe the selector of the collapse dynamics and the induced discrete-time evolution; in Section 4 we study reversibility for arbitrary (physically admissible) realization maps; Section 5 draws some conclusions; the Appendix is devoted to the proof of Theorem 4.5.
@@ 4 Reversibility
The following result is key to our reasoning. For the proof, see the Appendix. The proof is a refinement of that of Theorem 3.9 in [11]. To make the present work independent, we cover it in detail without relying on the proof of the weaker result.
@@
Observe now that an infinite strong chain-transitive subset $S$ is necessarily uncountable. Indeed, assume by contradiction that $S$ is countable. Since $S$ is closed, it is a compact metric space. Any countable compact metric space contains an isolated point $v]\in S$. By invariance, $T([v])\in S$, and since $[v]$ is isolated, chain-transitivity implies that $[v]$ is periodic, contradicting the absence of periodic orbits. Hence $|S|>\aleph_{0}$. ∎
@@
5 Conclusions
It is perhaps useful to frame this result by analogy with the logical structure of some classical thermodynamics concepts. The second law provides a universal lower bound on irreversibility, while Carnot’s theorem is an existence result showing that this bound can in fact be saturated: it identifies ideal cycles that achieve reversibility under suitable structural constraints. Landauer’s principle plays a similar role to the second law, including information in the entropy balance, furnishing a universal lower bound on dissipation for logically irreversible operations. Our result stands in a similar relation to Landauer’s principle as Carnot’s theorem does to the second law: it is an existence statement identifying conditions (compactness) under which the bound is asymptotically saturated. Observe, however, that Carnot’s theorem is an inherently stronger result, within classical thermodynamics, than ours within quantum reversibility theory. Indeed, Carnot’s theorem provides a universal criterion, whereas our result only identifies a suitable subsystem showing quasi-reversibility.
Declarations
The authors declare that they have no conflicts of interest.
The manuscript has no associated data.
Appendix
In this Appendix, we address the proof of Theorem 4.5. As stated in the Introduction, we aim to provide an argument that does not rely on the Axiom of Choice, so that the construction makes the link between the chain-transitive subsystem and the dynamics transparent.
We start recalling the following result that shows that in a compact dynamical system, without any regularity assumption on the map, there always exists a generalized recurrent point for the system, that is, a point $x\in X$ such that $x\in GR(f):=\bigcap_{d\in M}\mathcal{SCR}_{d}(f)$ where $M$ is the set of the metrics compatible with the topology of $X$.
@@ 5 Conclusions
The following lemma is a direct consequence of [11, Lemma 3.6 and Remark 3.8] and Remark 5.2.
@@
The proof of this lemma is a consequence of the construction in Lemma 5.1, and by using [11, Remark 3.8] and Remark 5.2, it is analogous to the proof of [11, Lemma 3.6]. In particular, from Lemma 5.3, we can say that for every $x\in X$, there exists a least ordinal $\alpha>0$ such that $x_{\alpha}\in S_{\alpha}(x)$ and thus $x_{\alpha}=x_{\beta}$ for some $\beta<\alpha$. Moreover, the set $S_{\beta,\alpha}(x)$ is an invariant and internally $\mathcal{SC}_{d}$-transitive subset of $X$.
@@
Now we are ready to prove Theorem 4.5, that for consistency we state here in abstract topological dynamical terms.
@@
We need to show that there is a closed, invariant, and internally $\mathcal{SC}_{d}$-transitive subset of $X$. Suppose, towards a contradiction, that $(X,f)$ does not have a $\mathcal{SC}_{d}$-transitive subsystem. In particular, we can assume that every orbit is infinite. Let $x^{0}\in X$. By Lemma 5.1, there exists a least ordinal $\alpha_{0}>0$ such that $x^{0}_{\alpha_{0}}\in S_{\alpha_{0}}(x^{0})$, and so $x^{0}_{\alpha_{0}}=x^{0}_{\beta_{0}}$ for some $\beta_{0}<\alpha_{0}$. By Lemma 5.3 and Remark 5.2, the set $S_{\beta_{0},\alpha_{0}}(x^{0})$ is an invariant, internally $\mathcal{SC}_{d}$-transitive subset. Since $S_{\beta_{0},\alpha_{0}}(x^{0})$ cannot be closed, there exists some $y\in(S_{\beta_{0},\alpha_{0}}(x^{0}))^{\prime}\setminus S_{\beta_{0},\alpha_{0}}(x^{0})$. Let $x\in S_{\alpha_{0}}(x^{0})$ and $\epsilon>0$, we want to show that there exists a strong $(\epsilon,d)$-chain from $x$ to $y$ whose points belong to $S_{\alpha_{0}}(x^{0})\cup\{y\}$. Let $\beta_{0}\leq\eta<\alpha_{0}$ be such that $d(x_{\eta},y)<\epsilon/2$. By property (9) and Remark 5.2, it follows that there exists a strong $(\epsilon/2,d)$-chain $x=y_{0},y_{1},\ldots,y_{n-1},y_{n}=x_{\eta}$ from $x$ to $x_{\eta}$ whose points belong to $S_{\alpha_{0}}(x^{0})$. By the triangle inequality,
@@
Note that the previous proof does not use the Axiom of Choice, as the operation of selecting a point from closed sets in a compact metric space can be performed by means of a selection function that exists in ZF+DC ([16])

### v4

# A no-go theorem for irreversibility along single-branch collapse dynamics

@@ Abstract
We study finite dimensional quantum systems with arbitrary collapse events, establishing, under no-information-erasure conditions, a structural no-go for operational irreversibility along single branches of the collapse dynamics. More precisely, we prove that, for every physically admissible selector of the collapse dynamics, there exists a topologically closed, forward-invariant subset of the projective state space on which any two states can be connected with arbitrarily fine Fubini–Study precision and arbitrarily small integrated energetic cost. This shows that the preservation of information along a realized branch of outcomes guarantees islands of quasi-reversibility, while genuine irreversibility requires additional ingredients such as non-compactness or information erasure.
@@ 1.1 Reversibility in unitary versus collapse dynamics
The seemingly intractable irregularity that such state–wise selectors permit is usually sidestepped, for instance by averaging over outcomes or retaining stochasticity, where continuity and stability are recovered in measure or in expectation rather than pointwise for a deterministic selector–induced map [9, 3, 2, 1, 21]. Our approach is instead to explore what can be said, from the point of view of recurrence, assuming only 1. and 2. above, so in particular allowing 3. We fix thus a realization map (the chosen selector of the dynamics) assuming only the (physics-dictated) properties 1-2 above, and allowing the density property (3). In particular, we allow that the the induced evolution map $T$ be everywhere discontinuous.
@@ 1.2 The result
Once a realization map $D$ has been chosen, and an ambient Hamiltonian $H$ is fixed, with corresponding unitary group $U$, they determine a (very irregular) induced dynamical evolution law:
@@
acting on unit vectors $u]\in\mathbb{P}(\mathcal{H})$.
Landauer’s principle (see [18]) bounds the minimal heat cost of logically irreversible operations (such as erasure), and conversely permits, in principle, arbitrarily low dissipation for logically reversible transformations in the quasi–static limit. The choice of a selector for the dynamics can be interpreted as realizing the extreme, no-information-erasure limit of Landauer’s principle, in the sense highlighted by C.H. Bennett’s resolution of the Maxwell demon paradox [5, 6, 13]. In particular, we share the following position, expressed by Bennett (from [5]):
@@
Given two states $u],[v]\in\mathbb{P}(\mathcal{H})$, we can define the minimal forward cost of going from $[u]$ to $[v]$ under the dynamical law $T_{U,D}$ by
@@
Let $T_{U,D}$ be a dynamical law and $B\subseteq\mathbb{P}(\mathcal{H})$ be a $T_{U,D}$-invariant subset of the state space. We say that the dynamical system $(B,T_{U,D})$ exhibits an operational arrow of time if there exist states $u],[v]\in\mathbb{P}(\mathcal{H})$ such that
@@
Conversely, we say that the operational arrow of time does not arise (or that $B$ is operationally reversible) if
@@
We can summarize our main result informally as follows.
Main result (informal) In the no–erasure regime, compactness enforces quasi–reversibility: given any finite-dimensional quantum system and any realization of the collapse dynamics, there exists a nonempty, topologically closed and invariant (with respect to the realization-induced evolution map) subsystem $S$ that is operationally reversible, meaning that, in $S$, quasi-returns with arbitrarily small Fubini-Study error and arbitrarily small energetic cost are allowed between any pair of states.
Genuine irreversibility, thus, requires additional ingredients: information loss/coarse–graining, non-compact limits, or coupling to reservoirs. This establishes a structural no-go: under finite-dimensional, no-erasure conditions, collapse alone cannot generate an operational arrow of time along individual realized branches.
@@ 1.3 Where a naive diagonal grid argument fails
It may look tempting to try to get a (naive) heuristic argument proving our main result as follows: fix $\epsilon>0$, cover $\mathbb{P}(\mathcal{H})$ by finitely many $d_{FS}$–balls ($\epsilon$–cells”), and note that along any long enough forward segment some cell must be visited twice; this yields an $\epsilon$–loop. Pushing $\epsilon\to 0$ seems to promise finer and finer quasi-loops. However, this argument is intrinsically scale–wise: as $\epsilon$ shrinks, the repeating cell and the base point inside it drift with $\epsilon$. Of course, by compactness the chosen cells admit a convergent subsequence, but in the absence of continuity the induced dynamics at the limit point is decoupled from the finite $\epsilon$–loops: $T$ at the limit may bear no relation to the images seen along the approximants. Consequently, the grid argument does not deliver a chain–recurrent point, let alone a topologically closed, invariant, internally strongly chain–transitive subsystem (which, as we will see, is the desired result); nothing in this argument enforces nesting across scales of the pseudo–orbits.
@@
1.4 Constructive Result over Non-Constructive Existence
At the opposite extreme of the naive attempt described above, one could obtain an abstract pure existence result on the existence of a chain-transitive subsystem by using Zorn’s lemma on topologically closed, invariant subsets ordered by inclusion. While this method does provide an existence result for a closed, invariant and chain transitive subset of the state space, it acts as a non-constructive “black box” that obscures the very operationality we seek to characterize.
Our aim, instead, is to give a constructive procedure, which is the only one, in our opinion, that works well with the operational character of the employed notion of reversibility. Therefore, not relying on the Axiom of Choice, we implement (in Section 5) a transfinite argument that enforces both spatial nesting and dynamical coherence across scales. This procedure yields a strongly chain–recurrent point, and even a topologically closed TT–invariant subset that is internally $\mathcal{SC}_{d_{FS}}$–transitive (see Sections 2 and 3 for the relevant definitions). Even more importantly, the two basic operations considered in the procedure, that is taking images of the map and passing to an accumulation point of an orbit, clearly have finite-precision analogues. The transfinite induction serves thus as a rigorous proxy for an arbitrary-precision search protocol, which Zorn’s lemma (or in general arguments intersecting a descending chain of suitable subsets, even if relying on weaker forms of choice) cannot provide.
The paper is organized as follows: in Section 2 we introduce basic concepts and notation for the unitary evolution of the system and the abstract topological dynamics concepts that will be used later; in Section 3 we describe the selector of the collapse dynamics and the induced discrete-time evolution; in Section 4 we study reversibility for arbitrary (physically admissible) realization maps; Section 5 is devoted to the proof of Theorem 4.5; Section 6 draws some conclusions.
@@ 4 Reversibility
The following result is key to our reasoning. For the proof, see Section 5.
@@
Observe now that an infinite, strong chain-transitive subset $S$ is necessarily uncountable. Indeed, assume by contradiction that $S$ is countable. Since $S$ is closed, it is a compact metric space. Any countable compact metric space contains an isolated point $v]\in S$. By invariance, $T([v])\in S$, and since $[v]$ is isolated, chain-transitivity implies that $[v]$ is periodic, contradicting the absence of periodic orbits. Hence $|S|>\aleph_{0}$. ∎
@@
5 Proof of Theorem 4.5
In this section, we address the proof of Theorem 4.5. This result is a strengthening of Theorem 3.9 in [11]. To make the present work independent, we prove it without relying on the proof of the weaker result.
As stated in the Introduction, we aim to provide an argument that does not rely on the Axiom of Choice, so that the construction makes the link between the chain-transitive subsystem and the dynamics transparent.
We start recalling the following result that shows that in a compact dynamical system $(X,f)$, without any regularity assumption on the map $f$, there always exists a generalized recurrent point for the system, that is, a point $x\in X$ such that $x\in GR(f):=\bigcap_{d\in M}\mathcal{SCR}_{d}(f)$ where $M$ is the set of the metrics compatible with the topology of $X$.
@@
The following lemma is a direct consequence of [11, Lemma 3.6] and Remark 5.2.
@@
The proof of this lemma follows by the construction in Lemma 5.1, and by using Remark 5.2, it is analogous to the proof of [11, Lemma 3.6]. Indeed, by Lemma 5.1, for every $x\in X$, there exists a least ordinal $\alpha>0$ such that $x_{\alpha}\in S_{\alpha}(x)$ and thus $x_{\alpha}=x_{\beta}$ for some $\beta<\alpha$. In the proof of [11, Lemma 3.6], it is shown that the set $S_{\beta,\alpha}(x)$ is invariant, and by Remark 5.2, $S_{\beta,\alpha}(x)$ is an internally $\mathcal{SC}_{d}$-transitive subset of $X$.
@@
We are now ready to prove Theorem 4.5, which we state here in abstract topological dynamical terms for consistency.
@@
We need to show that there is a closed, invariant, and internally $\mathcal{SC}_{d}$-transitive subset of $X$. Suppose, towards a contradiction, that $(X,f)$ does not have a $\mathcal{SC}_{d}$-transitive subsystem. In particular, we can assume that every orbit is infinite. Let $x^{0}\in X$. By the last step of the proof of Lemma 5.1, there exists a least ordinal $\alpha_{0}>0$ such that $x^{0}_{\alpha_{0}}\in S_{\alpha_{0}}(x^{0})$, and so $x^{0}_{\alpha_{0}}=x^{0}_{\beta_{0}}$ for some $\beta_{0}<\alpha_{0}$. By Lemma 5.3 and Remark 5.2, the set $S_{\beta_{0},\alpha_{0}}(x^{0})$ is an invariant, internally $\mathcal{SC}_{d}$-transitive subset. Since $S_{\beta_{0},\alpha_{0}}(x^{0})$ cannot be closed, there exists some $y\in(S_{\beta_{0},\alpha_{0}}(x^{0}))^{\prime}\setminus S_{\beta_{0},\alpha_{0}}(x^{0})$. Let $x\in S_{\alpha_{0}}(x^{0})$ and $\epsilon>0$, we want to show that there exists a strong $(\epsilon,d)$-chain from $x$ to $y$ whose points belong to $S_{\alpha_{0}}(x^{0})\cup\{y\}$. Let $\beta_{0}\leq\eta<\alpha_{0}$ be such that $d(x_{\eta},y)<\epsilon/2$. By property (9) and Remark 5.2, it follows that there exists a strong $(\epsilon/2,d)$-chain $x=y_{0},y_{1},\ldots,y_{n-1},y_{n}=x_{\eta}$ from $x$ to $x_{\eta}$ whose points belong to $S_{\alpha_{0}}(x^{0})$. By the triangle inequality,
@@
Note that the previous proof does not use the Axiom of Choice, as the operation of selecting a point from closed sets in a compact metric space can be performed by means of a selection function that exists in ZF+DC ([16]).
6 Conclusions
It is perhaps useful to frame this result by analogy with the logical structure of some classical thermodynamics concepts. The second law provides a universal lower bound on irreversibility, while Carnot’s theorem is an existence result showing that this bound can in fact be saturated: it identifies ideal cycles that achieve reversibility under suitable structural constraints. Landauer’s principle plays a similar role to the second law, including information in the entropy balance, furnishing a universal lower bound on dissipation for logically irreversible operations. Our result stands in a similar relation to Landauer’s principle as Carnot’s theorem does to the second law: it is an existence statement identifying conditions (compactness) under which the bound is asymptotically saturated.
Declarations
The authors declare that they have no conflicts of interest.
The manuscript has no associated data.